OpenStudy (anonymous):

write y=3x^2-6x+4 in vertex form. A. y=3(x-1)^2+1 B. y=3(x-1)^2+3 C. y=3(x-1)^2+5 D. y=3(x-1)^2+7

OpenStudy (anonymous):

Vertex form: y=a(x-h)^2+k y=3(x^2-2x)+4 y=3(x-2x+1)+4-3 y=3(x-1)^2+1 Your answer is a.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome!